The average price of 350 cellphones is ₱13,500 with the sample standard deviation of ₱750.

a. Find the 99% confidence interval of the true average price of the cellphones.

b. Find the length of the confidence interval.


1
Expert's answer
2021-05-25T17:44:17-0400

Part (1)


The Confidence Interval formula is



"\bar{X} ± Z \frac{s}{\sqrt{n}}"




Where:

  • X is the mean=13500
  • Z is the Z-value from the table below=2.576
  • s is the standard deviation=750
  • n is the number of observations=350


Note:-

Z =2.576 for 99% confidence.


Put all value in formula and we get


"13500 ± 2.576 \frac{750}{\sqrt{350}}"


Now solve ,we get

Confidence interval=(13500 ± 103)

=(13397 to 13603)

______________________________________

Part (2)


length of the confidence interval

= Upper value of interval - lower value of interval

=13603-13397

=206